OpenStudy (anonymous):

Which statement is true about the slope of the graph of {y = b^x}, where x and b are both greater than 1? a. Its slope decreases as x increases. b. Its slope varies between 1 and -1. c. Its slope increases as x increases.

OpenStudy (nincompoop):

well how about you do a little exposition

OpenStudy (mathmale):

Why not choose acceptable values for x and b and then actually graph the function y=b^x? Doing so would help you determine which of the three given answers is the best one. Try b = 2 (nice and simple). Then y=2^x. Let x = 1, 2, 3, 4. that should be enough. Graph those four points. Which of the three possible answers is the best one?

OpenStudy (mathmale):